OpenStudy (anonymous):

The position of a particle on the x-axis at time t, t > 0, is s(t) = e^t with t measured in seconds and s(t) measured in feet. What is the average velocity of the particle for 0 ≤ t ≤ In3? a. 3 b. 2/ ln 3 c. 3/ ln 3 d. 3/ 2

OpenStudy (jfraser):

average velocity is the rate of change in position over time. find the final position of the particle at t=3s

OpenStudy (anonymous):

@JFraser i still dont understand this

OpenStudy (jfraser):

find the position of the particle when t = 3s

OpenStudy (jfraser):

use the equation given \[s(t) = e^t\]to find the position when t = 3

OpenStudy (anonymous):

I first must find the derivative of s(t) = s'(t) = v(t) right?

OpenStudy (jfraser):

if you're looking for the \(average\) velocity, then no you don't need to find the derivative

OpenStudy (jfraser):

you're going to use the data given to find the \(average\) rate of change over the given period of time, not the \(instantaneous\) velocity at a particular point

OpenStudy (anonymous):

e^3?

OpenStudy (anonymous):

@JFraser

OpenStudy (jfraser):

that's the change in position of the particle after 3 seconds, \(e^3\) meters away from its starting point

OpenStudy (jfraser):

do the same thing, but instead of 3 seconds, do \(ln(3)\) seconds

OpenStudy (jfraser):

I just realized I misread the problem. It should have been \(ln(3)\) all along, not 3

OpenStudy (anonymous):

e^ln 3? @JFraser

OpenStudy (jfraser):

yes

OpenStudy (anonymous):

now what do i do?

OpenStudy (jfraser):

what's the value of \(e^{(ln(3))}\)

OpenStudy (anonymous):

3 because e and ln cancel each other out? @JFraser

OpenStudy (jfraser):

they are inverse functions, yes

OpenStudy (jfraser):

so now you have the position after \(ln(3)\) seconds, 3 meters away

OpenStudy (anonymous):

3 /ln 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

velocity = position/ distance right so 3 / ln 3. ln 3= distance

OpenStudy (jfraser):

average velocity is the rate of change of position over time. Now you have both of those values, the change in position is 3 meters, and the time is \(ln(3)\) seconds

OpenStudy (anonymous):

so 3 / ln 3?

OpenStudy (anonymous):

@JFraser

OpenStudy (jfraser):

that's the one

OpenStudy (anonymous):

Okay, thanks so much for the needed help! @JFraser

OpenStudy (anonymous):

This answer is wrong @JFraser

OpenStudy (anonymous):

I just got it wrong

OpenStudy (jfraser):

if the question's asking for the average velocity, that's what it is.

OpenStudy (jfraser):

I forgot that \(e^0\) isn't zero, it's one. So the change in position isn't 3 feet, it's 2 feet

OpenStudy (anonymous):

2 ln /3

OpenStudy (anonymous):

right? @JFraser

OpenStudy (anonymous):

hello? @JFraser

OpenStudy (anonymous):

@eliassaab can you help me?

OpenStudy (anonymous):

I have to ride the bus back to campus and leave. I won't be back online until 4:00ish

OpenStudy (anonymous):

*

OpenStudy (anonymous):

\[ \frac {s(\ln 3)-s(0)}{\ln (3)-0}= \frac{3-1}{\ln 3}=\frac 2 {\ln 3} \]

OpenStudy (anonymous):

The answer is b

OpenStudy (anonymous):

Thanks @eliassaab

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

You can practice problems like this on my site http://saab.org